# Computable conditions for order-2 $CP$ symmetry in NHDM potentials

### v1

@@ 1 Introduction
for some matrix $V\in\mathsf{SU}(N)$. In addition, a $CP$ symmetry need not be of order 2, with $CP^{2}$ being the identity, but may be of higher order $p=2q$, with $p>2$ the smallest integer such that $CP^{p}$ is the identity instead. In contrast to higher-order $CP$ symmetries, $CP2$ is equivalent to the existence of a basis where all the parameters are real [5]. In other words, $CP2$ is equivalent to the canonical $CP$ in some basis. NHDMs with CPs of higher order than 2 often generate $CP2$ as an accidental symmetry, and $CP2$ was long considered to be the only $CP$. However, while in the 2HDM $CP2$ is the only possible $CP$, a 3HDM with an order-4 $CP$ symmetry ($CP4$) and no other symmetries was identified and studied in [6, 7]. Even higher order $CP$s than $CP4$, with no accidental symmetries, were constructed and examined in [8].
@@
Throughout this paper we allow $N$ to be arbitrarily large, since our method in principle applies to any number of doublets, although its computational cost increases with $N$. While the 2HDM and 3HDM are currently the most relevant for phenomenology, models with more doublets have received some attention. 4HDMs were studied in e.g. [15, 16] and [17]. In the latter article, one doublet couples to quarks and three doublets couple to charged leptons, allowing for flavor changing neutral currents in the leptonic sector, but not in the quark sector. 5HDMs in the context of higher order $CP$s were scrutinized in [8]. A 6HDM for Dark Matter was examined in [18], and Grand Unified Theories with eight and nine Higgs doublets were studied in [19] and [20], respectively. Moreover, in the ”Private Higgs” extension of the SM each charged fermion acquire mass from its own Higgs doublet, through $\mathcal{O}(1)$ Yukawa couplings, and is hence another example of a model with $N=9$ Higgs doublets [21, 22]. The analysis of such models may be facilitated by the general algorithm for $CP2$ detection presented here.
@@ 2 Formalism
where the bilinears $K_{\alpha}$, $\alpha=0...N^{2}-1$ are given in terms of the generalized Gell-Mann matrices $\lambda_{a}$
$K_{0}=\Phi_{i}\Phi_{i}\;,\quad K_{a}=\Phi_{i}^{\dagger}(\lambda_{a})_{ij}\Phi_{j}.$ (2.2)
@@
$\lambda_{a}^{T}=-\lambda_{a}\quad\text{for}\quad a=1,...,k\equiv\frac{N(N-1)}{2}.$ (2.8)
As we will see in Sect. 2.1, the fact that this subset is equivalent to the defining representation of $\mathfrak{so}(N)$ can be used to derive simple necessary and sufficient conditions for $CP2$ symmetry in NHDMs.
@@ 2.1 Covariant framework for detecting $CP2$
where $f_{ijk}$ are the structure constants of $\mathfrak{su}(N)$ The isomorphism is then given by the map
@@
2.1.1 Necessary and sufficient condition for $CP2$ symmetry
@@
  • $k$ of $\Lambda$’s eigenvectors, $\{v_{a}\}_{a=1}^{k}$, form a basis for the defining representation of ${\mathfrak{so}}(N)$
  • $L\cdot v_{a}=M\cdot v_{a}=0,\quad\forall a\in\{1,...,k\}$.
@@
is block diagonal with $B_{N}$ and $A_{N}$ arbitrary symmetric matrices of dimension $k\times k$ and $N^{2}-1-k\times N^{2}-1-k$, respectively. In this basis, it is evident that $k$ of $\Lambda$’s eigenvectors span $E_{A}$, and therefore the image of this set by the isomorphism (2.11) is a basis for the defining representation of ${\mathfrak{so}}(N)=\text{Span}(\lambda_{1},...,\lambda_{k})$. Denote this subset of eigenvectors by $\{t_{a}\}_{a=1}^{k}$, it follows that
@@
$L\cdot t_{a}=M\cdot t_{a}=0,\quad\forall a\in\{1,...,k\}.$ (2.21)
@@
$UV_{a}U^{\dagger}\in\text{Span}(\lambda_{1},...,\lambda_{k}),\quad\forall a\in\{1,...,k\}.$ (2.22)
@@
$R(U)v_{a}\in E_{A},\quad\forall a\in\{1,...,k\}$ (2.23)
which implies these $k$ eigenvectors span $E_{A}$ and that the remaining ones $\{v_{a}\}_{a=k+1}^{N^{2}-1-k}$ span $E_{S}$. Hence, writing $\Lambda$ in terms of its eigenvectors $v_{a}$ and eigenvalues $\alpha_{a}$ using its spectral decomposition, it follows that
@@ 2.2 Identifying Lie algebras and representations
Applying the characterization of $CP2$ derived in the previous section requires identifying Lie algebras and their representations. Therefore, for the purpose of making the present paper self-contained, we now give a brief reminder of Lie algebra theory focusing on the classification of semi-simple Lie algebras and their representations. This presentation is not meant to be exhaustive but simply to introduce characteristics of Lie algebras and how they can be computed in practice. For more complete expositions of Lie algebra and representation theory see e.g. [25, 26, 27].
@@
while the $d-r$ eigenvalue tuples $\alpha=(\alpha_{1},...,\alpha_{r})$, thought of as vectors of $\mathbb{R}^{r}$, form the so-called root system of $\mathfrak{g}$. It was shown by Dynkin that semi-simple Lie algebras can be classified according to their root system [28].
Thus an unknown semi-simple Lie algebra can be identified if one can compute a Cartan subalgebra for it. This can be done by calculating the nullspace of an adjoint matrix ${\text{ad}}_{X}$ where $X$ is, by definition, any regular element of $\mathfrak{g}$ [30, 29]. For ${\mathfrak{su}}(N)$ and its subalgebras [^2] If $\mathfrak{g^{\prime}}\subset\mathfrak{g}$ and $x$ is regular in $\mathfrak{g}$ then $x$ is regular in $\mathfrak{g^{\prime}}$ [30]., an element is regular if all its eigenvalues are distinct. Thus a Cartan subalgebra can be computed from a generic element e.g. randomly sampled.
The dimension of the nullspace of ${\text{ad}}_{X}$ then gives the rank of $\mathfrak{g}$, and the nullvectors $h_{i}$ provide a basis $H_{i}$ for a Cartan subalgebra. One can then simultaneously diagonalize the matrices ${\text{ad}}_{H_{i}}$ and find the root system. Fig 1 shows how ${\mathfrak{so}}(7)$ and $\mathfrak{sp}(6)$, which have the same dimension and rank, differ by their root system.
Once the root system $R$ has been found and an ordered set of positive simple roots $\{\beta_{1},...,\beta_{r}\}\subset R$ has been chosen [25], the Lie algebra representation at hand can be identified by computing its highest weight $\Upsilon$ which, for an n-dimensional irreducible representation, is a vector of $\mathbb{R}^{r}$ characterizing that representation [26]. In the case of a reducible representation there will be one highest weight per irreducible component. To each highest weight corresponds a simultaneous nullvector $v_{0}$ of the positive simple roots
$E_{\beta_{i}}v_{0}\equiv 0\,,\quad\forall i\in\{1,...,r\}.$ (2.28)
The components of the highest weight $\Upsilon=(a_{1},...,a_{r})$ in the basis of so-called fundamental weights $\{\omega_{i}\}_{i=1}^{r}$ [26], sometimes called Dynkin labels, are the smallest integers satisfying
$E_{-\beta_{i}}^{1+a_{i}}~{}v_{0}=0\,,\quad i\in\{1,...,r\}.$ (2.29)
@@ 2.2.1 ${\mathfrak{so}}(N)$ subalgebras of ${\mathfrak{su}}(N)$
In the case $\mathfrak{so}(2r)_{\mathbb{C}}$ with $r\geq 4$, the dimension of the defining representation, dim$(\Gamma_{\omega_{1}})=2r<\Gamma_{\omega_{k}}$ for $21<k\leq r-$ 2 by (2.2.1), while for the cases $k>r-2$, dim$(\Gamma_{\omega_{1}})=2r<2^{r-1}$ when $r>4$, so the defining representation has the uniquely least dimension among the fundamental representations, for $r>4$. And again, due to (2.33), the defining representation of $\mathfrak{so}(2r)_{\mathbb{C}}$ become the unique irreducible representation with dimension $2r$ for $r>4$. Note that for $r=4$ (i.e. $N=8$), (2.2.1) gives dim$(\Gamma_{\omega_{1}})=$dim$(\Gamma_{\omega_{3}})=$ dim$(\Gamma_{\omega_{4}})=8$, but this is one of the two cases the Proposition is not valid.
@@
The ”exceptional” subalgebras of Proposition 2 are consistent with the low $N$ subalgebra tables of [31] and [35]. Ref. [35] lists complex subalgebras of complex, simple algebras, but for ${\mathfrak{su}}(N)$, every complex subalgebra of the complexified algebra ${\mathfrak{su}}(N)_{\mathbb{C}}$ will correspond to a real subalgebra of the compact, real algebra ${\mathfrak{su}}(N)$, and vice versa. The latter direction holds for all algebras, while real ${\mathfrak{su}}(N)$ have real subalgebras in 1-1 correspondence with the complex subalgebras of ${\mathfrak{su}}(N)_{\mathbb{C}}$. The reason is a complex subalgebra $\mathfrak{h}_{\mathbb{C}}\subset{\mathfrak{su}}(N)_{\mathbb{C}}$ also is a faithful, complex representation of the subalgebra. And since there is a 1-1 correspondence between complex representations of real and complex variants of the algebras, there will also be a corresponding complex representation of the real, compact form $\mathfrak{h}$ of $\mathfrak{h}_{\mathbb{C}}$. This representation will exist in the real algebra ${\mathfrak{su}}(N)$, since every representation of a compact algebra is equivalent with a Hermitian representation [34], i.e. it is found among the Hermitian matrices of ${\mathfrak{su}}(N)$.
@@ 3.2 Detecting the defining representation of ${\mathfrak{so}}(N)$ in ${\mathfrak{su}}(N)$
The strategy is to first determine if a subset of $k$ eigenvectors of $\Lambda$ closes under the F-product i.e. forms a $k$-dimensional subalgebra. If such a subalgebra exists, we must verify whether it is the ${\mathfrak{so}}(N)$ algebra or some other $k$-dimensional subalgebra, since e.g. both ${\mathfrak{so}}(5)$ and ${\mathfrak{su}}(3)\oplus\mathfrak{u}(1)\oplus\mathfrak{u}(1)$ are 10-dimensional subalgebras of ${\mathfrak{su}}(5)$. For even $N=2r$, computing the rank of the unknown algebra using the method described in 2.2 is enough to unambiguously identify ${\mathfrak{so}}(2r)$, as it is the only [^3] We have checked this up to $N=22$ (rank 11) by exhausting all the possible semi-simple Lie algebras of rank $r$ and dimension $r(2r-1)$ and checking against the ${\mathfrak{su}}(N)$ subalgebra tables of [31]. For $N\geq 24$, there may be subalgebras of same dimension and rank as ${\mathfrak{so}}(N)$ and one has to look at the root systems. subalgebra of ${\mathfrak{su}}(2r)$ with dimension $k$ and rank $r$. For odd $N=2r+1$, ${\mathfrak{su}}(2r+1)$ always has, in addition to ${\mathfrak{so}}(2r+1)$, at least an $\mathfrak{sp}(2r)$ subalgebra which has the same dimension and rank. Thus, beyond $r=1$ and $r=2$ where one has the isomorphisms ${\mathfrak{so}}(3)\cong\mathfrak{sp}(2)$ and ${\mathfrak{so}}(5)\cong\mathfrak{sp}(4)$, the root system of the unknown algebra must be computed in order to establish that it is ${\mathfrak{so}}(2r+1)$.
@@ 3.2.1 Testing for a subalgebra
$Z_{ijk}=F^{(v_{i},v_{j})}\cdot v_{k}=\frac{-i}{4}{\rm tr}\big{(}[V_{i},V_{j}]V_{k}\big{)},$ (3.1)
and the closure of a subalgebra generated by a subset $\{v_{a}\}_{a\in I}$, $I\subset\{1,...,N^{2}-1\}$ means that

### v2

@@ 1 Introduction
for some matrix $V\in\mathsf{U}(N)$ [5]. In addition, a $CP$ symmetry need not be of order 2, with $CP^{2}$ being the identity, but may be of higher order $p=2q$, with $p>2$ the smallest integer such that $CP^{p}$ is the identity instead. In contrast to higher-order $CP$ symmetries, $CP2$ is equivalent to the existence of a basis where all the parameters are real [6]. In other words, $CP2$ is equivalent to the canonical $CP$ in some basis. NHDMs with CPs of higher order than 2 often generate $CP2$ as an accidental symmetry, and $CP2$ was long considered to be the only $CP$. However, while in the 2HDM $CP2$ is the only possible $CP$, a 3HDM with an order-4 $CP$ symmetry ($CP4$) and no other symmetries was identified and studied in [7, 8]. Even higher order $CP$s than $CP4$, with no accidental symmetries, were constructed and examined in [9].
@@
Throughout this paper we allow $N$ to be arbitrarily large, since our method in principle applies to any number of doublets, although its computational cost increases with $N$. While the 2HDM and 3HDM are currently the most relevant for phenomenology, models with more doublets have received some attention. 4HDMs were studied in e.g. [16, 17] and [18]. In the latter article, one doublet couples to quarks and three doublets couple to charged leptons, allowing for flavor changing neutral currents in the leptonic sector, but not in the quark sector. 5HDMs in the context of higher order $CP$s were scrutinized in [9]. A 6HDM for Dark Matter was examined in [19], and Grand Unified Theories with eight and nine Higgs doublets were studied in [20] and [21], respectively. Moreover, in the “Private Higgs” extension of the SM each charged fermion acquire mass from its own Higgs doublet, through $\mathcal{O}(1)$ Yukawa couplings, and is hence another example of a model with $N=9$ Higgs doublets [22, 23]. The analysis of such models may be facilitated by the general algorithm for $CP2$ detection presented here.
@@ 2 Formalism
where the bilinears $K_{\alpha}$, $\alpha=0,\ldots,N^{2}-1$ are given in terms of the generalized Gell-Mann matrices $\lambda_{a}$
$K_{0}=\Phi_{i}^{\dagger}\Phi_{i}\,,\quad K_{a}=\Phi_{i}^{\dagger}(\lambda_{a})_{ij}\Phi_{j}.$ (2.2)
@@
$\lambda_{a}^{T}=-\lambda_{a}\quad\text{for}\quad a=1,\ldots,k\equiv\frac{N(N-1)}{2}.$ (2.8)
As we will see in section 2.1, the fact that this subset is equivalent to the defining representation of $\mathfrak{so}(N)$ can be used to derive simple necessary and sufficient conditions for $CP2$ symmetry in NHDMs.
@@ 2.1 Covariant framework for detecting $CP2$
where $f_{ijk}$ are the structure constants of $\mathfrak{su}(N)$ in the Gell-Mann basis. The isomorphism is then given by the map
@@
2.1.1 Necessary and sufficient conditions for $CP2$ symmetry
@@
  • $k=\frac{N(N-1)}{2}$ of $\Lambda$’s eigenvectors, $\{v_{a}\}_{a=1}^{k}$, form a basis for the defining representation of ${\mathfrak{so}}(N)$
  • $L\cdot v_{a}=M\cdot v_{a}=0,\quad\forall a\in\{1,\ldots,k\}$.
@@
is block diagonal with $B_{N}$ and $A_{N}$ arbitrary symmetric matrices of dimension $k\times k$ and $N^{2}-1-k\times N^{2}-1-k$, respectively. In this basis, it is evident that $k$ of $\Lambda$’s eigenvectors span $E_{A}$, and therefore the image of this set by the isomorphism (2.11) is a basis for the defining representation of ${\mathfrak{so}}(N)=\text{Span}(\lambda_{1},\ldots,\lambda_{k})$. Denote this subset of eigenvectors by $\{t_{a}\}_{a=1}^{k}$, it follows that
@@
$L\cdot t_{a}=M\cdot t_{a}=0,\quad\forall a\in\{1,\ldots,k\},$ (2.21)
@@
$UV_{a}U^{\dagger}\in\text{Span}(\lambda_{1},\ldots,\lambda_{k}),\quad\forall a\in\{1,\ldots,k\}.$ (2.22)
@@
$R(U)v_{a}\in E_{A},\quad\forall a\in\{1,\ldots,k\}$ (2.23)
which implies these $k$ eigenvectors span $E_{A}$ and that the remaining ones $\{R(U)v_{a}\}_{a=k+1}^{N^{2}-1}$ span $E_{S}$. Hence writing $\Lambda$ in terms of its eigenvectors $v_{a}$ and eigenvalues $\alpha_{a}$ using its spectral decomposition it follows that
@@ 2.2 Identifying Lie algebras and representations
Applying the characterization of $CP2$ derived in the previous section requires identifying Lie algebras and their representations. Therefore, for the purpose of making the present paper self-contained, we now give a brief reminder of Lie algebra theory focusing on the classification of semisimple Lie algebras and their representations. This presentation is not meant to be exhaustive but simply to introduce characteristics of Lie algebras and how they can be computed in practice. For more complete expositions of Lie algebra and representation theory see e.g. [26, 27, 28].
@@
while the $d-r$ eigenvalue tuples $\alpha=(\alpha_{1},\ldots,\alpha_{r})$, thought of as vectors of $\mathbb{R}^{r}$, form the so-called root system of $\mathfrak{g}$. It was shown by Dynkin that semisimple Lie algebras can be classified according to their root system [29].
Thus an unknown semisimple Lie algebra can be identified if one can compute a Cartan subalgebra for it. This can be done by calculating the nullspace of an adjoint matrix ${\text{ad}}_{X}$ where $X$ is, by definition, any regular element of $\mathfrak{g}$ [30, 31]. For ${\mathfrak{su}}(N)$ and its subalgebras, [^2] If $\mathfrak{g^{\prime}}\subset\mathfrak{g}$ and $x$ is regular in $\mathfrak{g}$ then $x$ is regular in $\mathfrak{g^{\prime}}$ [30]. an element is regular if all its eigenvalues are distinct. Thus a Cartan subalgebra can be computed from a generic element e.g. randomly sampled.
The dimension of the nullspace of ${\text{ad}}_{X}$ then gives the rank of $\mathfrak{g}$, and the nullvectors $h_{i}$ provide a basis $H_{i}$ for a Cartan subalgebra. One can then simultaneously diagonalize the matrices ${\text{ad}}_{H_{i}}$ and find the root system. Figure 1 shows how ${\mathfrak{so}}(7)$ and $\mathfrak{sp}(6)$, which have the same dimension and rank, differ by their root system.
(a) ${\mathfrak{so}}(7)$
(b) $\mathfrak{sp}(6)$
Once the root system $R$ has been found and an ordered set of positive simple roots $\{\beta_{1},\ldots,\beta_{r}\}\subset R$ has been chosen [26], the Lie algebra representation at hand can be identified by computing its highest weight $\Upsilon$ which, for an n-dimensional irreducible representation, is a vector of $\mathbb{R}^{r}$ characterizing that representation [27]. In the case of a reducible representation there will be one highest weight per irreducible component. To each highest weight corresponds a simultaneous nullvector $v_{0}$ of the positive simple roots
$E_{\beta_{i}}v_{0}\equiv 0\,,\quad\forall i\in\{1,\ldots,r\}.$ (2.28)
The components of the highest weight $\Upsilon=(a_{1},\ldots,a_{r})$ in the basis of so-called fundamental weights $\{\omega_{i}\}_{i=1}^{r}$ [27], sometimes called Dynkin labels, are the smallest integers satisfying
$E_{-\beta_{i}}^{1+a_{i}}~{}v_{0}=0\,,\quad i\in\{1,\ldots,r\}.$ (2.29)
@@ 2.2.1 ${\mathfrak{so}}(N)$ subalgebras of ${\mathfrak{su}}(N)$
In the case $\mathfrak{so}(2r)_{\mathbb{C}}$ with $r\geq 4$, the dimension of the defining representation, dim$(\Gamma_{\omega_{1}})=2r<\Gamma_{\omega_{k}}$ for $21<k\leq r-$ 2 by (2.2.1), while for the cases $k>r-2$, dim$(\Gamma_{\omega_{1}})=2r<2^{r-1}$ when $r>4$, so the defining representation has the uniquely least dimension among the fundamental representations, for $r>4$. And again, due to (2.33), the defining representation of $\mathfrak{so}(2r)_{\mathbb{C}}$ becomes the unique irreducible representation with dimension $2r$ for $r>4$. Note that for $r=4$ (i.e. $N=8$), (2.2.1) gives dim$(\Gamma_{\omega_{1}})=$dim$(\Gamma_{\omega_{3}})=$ dim$(\Gamma_{\omega_{4}})=8$, but this is one of the two cases the Proposition is not valid.
@@
The “exceptional” subalgebras of Proposition 2 are consistent with the low $N$ subalgebra tables of [32] and [36]. Ref. [36] lists complex subalgebras of complex, simple algebras, but for ${\mathfrak{su}}(N)$, every semisimple complex subalgebra of the complexified algebra ${\mathfrak{su}}(N)_{\mathbb{C}}$ will correspond to a semisimple real subalgebra of the compact, real algebra ${\mathfrak{su}}(N)$, and vice versa. The latter direction holds for all algebras, while compact ${\mathfrak{su}}(N)$ have real semisimple subalgebras in 1-1 correspondence with the semisimple complex subalgebras of ${\mathfrak{su}}(N)_{\mathbb{C}}$. The reason is a complex subalgebra $\mathfrak{h}_{\mathbb{C}}\subset{\mathfrak{su}}(N)_{\mathbb{C}}$ also is a faithful, complex representation of the subalgebra. And since there is a 1-1 correspondence between complex representations of real and complex variants of the algebras, there will also be a corresponding complex representation of the real, compact form $\mathfrak{h}$ of $\mathfrak{h}_{\mathbb{C}}$. This representation will exist in the real algebra ${\mathfrak{su}}(N)$, since every representation of a compact algebra is equivalent with a Hermitian representation [35], i.e. it is found among the Hermitian matrices of ${\mathfrak{su}}(N)$.
@@ 3.2 Detecting the defining representation of ${\mathfrak{so}}(N)$ in ${\mathfrak{su}}(N)$
The strategy is to first determine if a subset of $k$ eigenvectors of $\Lambda$ closes under the F-product i.e. forms a $k$-dimensional subalgebra. If such a subalgebra exists, we must verify whether it is the ${\mathfrak{so}}(N)$ algebra or some other $k$-dimensional subalgebra, since e.g. both ${\mathfrak{so}}(5)$ and ${\mathfrak{su}}(3)\oplus\mathfrak{u}(1)\oplus\mathfrak{u}(1)$ are 10-dimensional subalgebras of ${\mathfrak{su}}(5)$. For even $N=2r$, computing the rank of the unknown algebra using the method described in 2.2 is enough to unambiguously identify ${\mathfrak{so}}(2r)$, as it is the only [^3] We have checked this up to $N=22$ (rank 11) by exhausting all the possible semisimple Lie algebras of rank $r$ and dimension $r(2r-1)$ and checking against the ${\mathfrak{su}}(N)$ subalgebra tables of [32]. For $N\geq 24$, there may be subalgebras of same dimension and rank as ${\mathfrak{so}}(N)$ and one has to look at the root systems. subalgebra of ${\mathfrak{su}}(2r)$ with dimension $k$ and rank $r$. For odd $N=2r+1$, ${\mathfrak{su}}(2r+1)$ always has, in addition to ${\mathfrak{so}}(2r+1)$, at least an $\mathfrak{sp}(2r)$ subalgebra which has the same dimension and rank. Thus, beyond $r=1$ and $r=2$ where one has the isomorphisms ${\mathfrak{so}}(3)\cong\mathfrak{sp}(2)$ and ${\mathfrak{so}}(5)\cong\mathfrak{sp}(4)$, the root system of the unknown algebra must be computed in order to establish that it is ${\mathfrak{so}}(2r+1)$.
@@ 3.2.1 Testing for a subalgebra
$Z_{ijk}\equiv F^{(v_{i},v_{j})}\cdot v_{k}=\frac{-i}{4}{\rm tr}\big{(}[V_{i},V_{j}]V_{k}\big{)},$ (3.1)
and the closure of a subalgebra generated by a subset $\{v_{a}\}_{a\in I}$, $I\subset\{1,\ldots,N^{2}-1\}$ means that